Write the equation of the line that passes through the piont (4,-9) with the slope 1/5 in piont-slope form
Found 2 solutions by Nate, checkley71:
Answer by Nate(3500)
Answer by checkley71(8403)

Y=mX+b OR Y=1/5X+b or the Y INTERCEPT IS -9-(4*1/5) OR -9-.8 OR -9.8 THUS THE EQUATION IS
Y=1/5X-9.2